Project cash flow and NPV. The managers of Classic Autos Incorporated plan to manufacture classic Thunderbirds (1957 replicas). The necessary foundry equipment will cost a total of $4,100,000 and will be depreciated using a five-year MACRS life,囲. Pro ected sales in annual units for the next five years are 290 per year. If the sales price is $28,000 per car, variable costs are $18,000 per car, and fixed costs are $1,300,000 annually, what is the annual operating cash flow f the tax rate is 30%? The equipment is sold for salvage for $550,000 at the end of year five. what is the after tax cash flow of the salvage? Net working capital increases by $575,000 at the beginning of the pro ect year 0 and is reduced back to its original level in the final year what is the incremental cash flow of the pro ect? Using a discount rate of 14% for the project, determine whether the project should be accepted or rejected according to the NPV decision model. First, what is the annual operating cash flow of the project for year 1? 1 Data Table $(Round to the nearest dollar.) MACRS Fixed Annual Expense Percentages by Recovery Class 5-Year 20.00% 32.00% 19.20% 11.52% 11.52% 5.76% 10-Year 10.00% 18.00% 14.40% 11 .52% 9.22% 737% 6.55% 6.55% 6.55% 6.55% 3.28% 3- Year 33.33% 44.45% 14.81% 7.41% 7-Year 14.29% 24.49% 17.49% 12.49% 8.93% 8.93% 8.93% 4.45% 10
Answer #1
Sales $                8,120,000 =290*28000
Less:
Costs $                6,520,000 =290*18000+1300000
Depreciation $                   820,000 =4100000*20%
EBT $                   780,000
Less: Tax payable @ 30% $                   234,000
Incremental earnings $                   546,000
Add: Depreciation $                   820,000
Operating cash flow $ 1,366,000